Jake Entertainment Corporation has three segments with revenue, operating income, and depreciation and amortization information (in millions) as follows: Operating Depreciation Segment Revenue Income and Amortization Film Theme Park Video Game Totals $5,000 1,000 500 $1,500 320 175 $525 112 53 $6,500 $1,995 $690 The segment with the highest EBITDA as a percent of revenue is

A. Video Game

B. The Theme Park and Video Game segments have equal EBITDA as a percent of revenue.
Answer #1
Film Theme park Video game
EBITDA 1500+525=2025 320+112 = 432 175+53 = 228
Revenue 5000 1000 500
EBITDA/revenue*100 40.5 43.2 45.6

So, Video game segment gives the highest EBITDA as a percent of revenue.

Please comment in case of any query regarding the solution